OpenStudy (anonymous):

According to the Americans with Disabilities Act, a ramp can rise no more than 1 ft for every 12 ft of horizontal distance. What is the maximum angle that a ramp can form with the ground?

OpenStudy (anonymous):

what is it?

OpenStudy (john_es):

Calculate the tangent, using that the horizontal distance is b=12, and vertical a=1, \[\tan\alpha=\frac{a}{b}\] Then use the result to calculate the angle, \[\alpha=\arctan(a/b)\]